Remarks on Self-Affine Tilings

Derek Hacon Nicolau C. Saldanha J. J. P. Veerman

Abstract

We study self-affine tilings of ℝ^{n} with special emphasis on the twodigit case. We prove that in this case the tile is connected and, if n ≤ 3, is a lattice-tile.